OpenStudy (anonymous):

Juice-O-Rama is a refreshment stand specializing in a fruit drink mixture of orange juice and prune juice. The orange juice costs $1.20 per liter and the prune juice costs $1.40 per liter. If they need 8 liters of the final drink at a cost of $1.24 per liter, how many liters of each juice will they need to make their specialty drink?

OpenStudy (anonymous):

Hi there. Did you just want an answer or do you want to learn the material?

OpenStudy (anonymous):

I want to learn. I dont understand.

OpenStudy (anonymous):

OK. So call 'x' the number of liters of orange juice, and 'y' the number of liters of prune juice'. We need to write two equations that relate X to Y. First, we need a total of 8 liters of juice. so, how can you write this in terms of X and Y?

OpenStudy (anonymous):

x+y=8

OpenStudy (anonymous):

yes!. nice. And, we want the total cost to be 1.24 per liter. So let's do this equation step by step. If it costs 1.24 per liter and we want 8 liters, what's the total cost?

OpenStudy (anonymous):

1.24(8) = 9.92

OpenStudy (anonymous):

yep. And, the total cost is equal to the sum of the cost of the orange juice and the cost of the prune juice. so if we have X liters of orange, and Y liters of prune, how much does the total cost, given the cost per liter of each juice in the problem?

OpenStudy (anonymous):

this is where i am lost.

OpenStudy (anonymous):

ok. How much is orange juice per liter?

OpenStudy (anonymous):

1.20

OpenStudy (anonymous):

and we have X liters of it. So in the final mixture, there is how much worth of orange juice in it?

OpenStudy (anonymous):

1.20x+1.40y=9.92?

OpenStudy (anonymous):

you got it my friend. that's our second equation. now solve that simultaneously with the first, x+y=8

OpenStudy (anonymous):

we have X liters of OJ, and Y liters of prune, and the total value of the 8 liters is 9.92.

OpenStudy (anonymous):

how do i solve it? (-1) x+y=8 -x-y=-8?

OpenStudy (anonymous):

for this one the substitution method is probably easier than the addition method, i.e. x+y=8 x = 8-y then substitute x=8-y into equation 2.

OpenStudy (anonymous):

x=6.4 y =1.6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that's what I got!

OpenStudy (anonymous):

Thank You so much. I just have one question. How do I know when to use the addition method vs. the substitution method.

OpenStudy (anonymous):

Honestly, whichever looks easier to you. Both methods are guaranteed to give you the same answer, but sometimes one has easier math than the other. I picked substitution method for that one because I could express x without a fraction. Fractions always make the math harder. If there's something like this: 2x + 8y = 3 2x - 3y = 7 Then I'll usually use the addition method because trying to isolate x or y in either of those equations would result in ugly fractions, whereas I can see easily that just reversing the sign of the second equation and adding gives me 11y = 10.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Oops, I mean 11y = -4.... ha. I'm not good at math ;-)